OpenStudy (ilovecake):

What is the value of r + s + t when r = 26, s = –12, and t = –8? A. 6 B. 22 C. 30 D. 46

OpenStudy (anonymous):

you need help

OpenStudy (ilovecake):

yep

TheSmartOne (thesmartone):

r+s+t= r=26 s=-12 t=-8

TheSmartOne (thesmartone):

Just plug in the numbers into there espective spots.

OpenStudy (anonymous):

6

OpenStudy (ilovecake):

SO I add all of those up?

TheSmartOne (thesmartone):

So it becomes 26+(-12)+(-8)

OpenStudy (anonymous):

yea that but put the negatives in parenthesis

TheSmartOne (thesmartone):

Can you solve it from there @Ilovecake

OpenStudy (ilovecake):

so answer is 6?

OpenStudy (anonymous):

yes 6

TheSmartOne (thesmartone):

Let me see

OpenStudy (anonymous):

yea

TheSmartOne (thesmartone):

26+(-12)+(-8) 26-12-8

TheSmartOne (thesmartone):

=14-8 =6

TheSmartOne (thesmartone):

yup exactly @Ilovecake

TheSmartOne (thesmartone):

yes @ilovecake you have to add them all up

OpenStudy (ilovecake):

Do you have time for one more?

TheSmartOne (thesmartone):

\[a+(-b)=a-b\]

TheSmartOne (thesmartone):

Sure @Ilovecake

TheSmartOne (thesmartone):

btw I like your username... lol

OpenStudy (ilovecake):

Thanks. Which statements are true? Choose all answers that are correct. A. –13 + 25 = 10 + (–22) B. 20 + (–6) > 31 + (–19) C. 8 + (–12) < –17 + 14 D. –11 + (–15) > 6 + (–29)

OpenStudy (anonymous):

i can't understand

OpenStudy (ilovecake):

Try refreshing the page

TheSmartOne (thesmartone):

What do you think the answer is @Ilovecake

OpenStudy (ilovecake):

i have no idea

OpenStudy (anonymous):

c is wrong

TheSmartOne (thesmartone):

Ok lets go one by one

OpenStudy (ilovecake):

brb

TheSmartOne (thesmartone):

A.–13 + 25 = 10 + (–22) Add -13+25=12 Then add 10+(-22)=-12 12 does not equal to -12

TheSmartOne (thesmartone):

So A is out

OpenStudy (anonymous):

d is out add -11+(-15)=-2 add 6+(-29)=-23 and -2 is not larger than -23